# RIFT: Entropy-Optimised Fractional Wavelet Constellations for Ideal Time–Frequency Estimation

### v2

@@ III-A Proposed Wavelet Transform
$\theta$ is the intended angle of the instantaneous angular frequency relative to the horizontal axis (time axis) in the T–F plane, and $\sigma$ is the intended standard deviation of the major axis of the corresponding Cohen’s class kernel. As explored previously in the overview (Section I-B), the employment of Short-time Fractional Fourier (or wavelet) Transforms is especially desirable in the context of time-varying signal curvature compared to their standard counterparts (for instance Short-Time Fourier Transforms or Wavelet Transforms), given their ability to align with signal trajectories and thus minimise cross term interference appropriately [11]. A demonstration of this phenomenon is provided in the SM, Section B. Employing the result of Eq. (14), the exact Cohen’s class smoothing kernel, $\bar{\Pi}_{\sigma,\theta}\left(\omega,\,t\right)$, resulting from the wavelet convolutional process can be obtained as:
@@
For the derivation, see the SM, Section C. Note that due to the skewed form of the kernel in Eq. (19), the intended angle and standard deviation of the principal axis of the kernels may not necessarily coincide with the actual values. The required mapping between the intended pair $(\sigma,\theta)$ (principal-axis angle and standard deviation) to the implemented parameters $(\sigma_{0},\kappa)$ is given by:
@@
where $K_{\sigma}\!\left(\kappa\right)$ is the axis switching function provided in Eq. (94) in the SM. Thus, $\kappa(\sigma,\theta)=\varphi^{-1}(\sigma,\theta)$, the inverse function with respect to $\theta$. Note that in the limiting case $\sigma\to\infty$, we obtain $\varphi(\sigma,\theta)\to\theta$ as expected. Equivalently, the mapping $(\sigma,\theta)\mapsto(\sigma_{0},\kappa)$ can be obtained by performing covariance matching with the intended Cohen’s class kernel (introduced in Eq. (25)); further details and derivations are given in the SM, Sections D1 and D2.
@@
A visualisation of the effect of the wavelet function on expected signal characteristics is provided in Section E of the SM.
@@ III-B Reconstructing the ITFR
See SM, Section F, for the full derivation. However, it is known that the solution for $\hat{\text{ITFR}}$ needs to be strictly non-negative from Eq. (33). Thus, we propose a nonnegativity-constrained solution to these normal equations (Eq. (114)) via Lucy–Richardson Deconvolution with total variation regularisation (LR-TV) [48]. TV regularisation, implemented via the multiplicative divergence term in Eq. (40) below, penalises the sum over T–F pixels of the Euclidean ($\ell_{2}$) norm of the local T–F gradient (isotropic TV). TV suppresses isolated pixel-scale speckle and deconvolution ringing while preserving edge-like ridges [48]. This standard LR-TV form [48] yields the following iterative update for the RIFT estimate of the ITFR, with respect to a spatially varying PSF, $\mathbf{\Psi}$:
@@
To efficiently compute the spatially varying Lucy–Richardson (LR) process, the overlap-save method can be employed [52, 53]. This approach involves performing parallel LR deconvolutions on $L\times H$ partitioned blocks of the 2D array. The results are then “stitched” together (with appropriate padding) to approximate the spatially varying process while maintaining computational efficiency. Full implementation details are provided in Section G in the SM. Likewise, a breakdown of the global RIFT process is provided in Algorithm 1.
@@ III-D Entropic Filtering for Cross-term Suppression
A complete derivation of this result is provided in the SM, Section H. It is recommended that $w_{u,v}$ is a bivariate Gaussian, with kernel variances selected to optimise performance for the desired application. For instance, for an instrumental or speech extract with known linearly spaced harmonics, the $\omega$ axis variance can be set high to enforce harmonic trajectory alignment in the presence of noise, otherwise, the kernel can be set to an isotropic Gaussian of variance greater than that of the derived variance of the isotropic Gaussian kernel in Eq. (45) to ensure valid entropic weightings.
@@ III-E Fundamental Properties of the RIFT
Note that, by construction, the RIFT output, $\widehat{I}(\omega,t)$, is positive ($\widehat{I}(\omega,t)\!\geq\!0$), time–shift and frequency–shift equivariant up to discretisation ($\widehat{I}_{x_{\tau}}(\omega,t)=\widehat{I}_{x}(\omega,t-\tau)$ for $x_{\tau}(t)=x(t-\tau)$ and $\widehat{I}_{x_{\nu}}(\omega,t)=\widehat{I}_{x}(\omega-\nu,t)$ for $x_{\nu}(t)=x(t)\,e^{j2\pi\nu t}$). See Section I in the SM for more details. It is also non-linear, specifically quadratic, due to the magnitude–squared of the CFWT coefficients (corresponding to energy). Likewise, by construction, the RIFT estimates the squared modulus of the ideal time–frequency representation (no phase information) and is thus non-invertible. Also, note that whilst the RIFT substantially suppresses cross-terms, it does not guarantee their elimination in all cases. Likewise, the target is a cross-term-free, WVD-resolution energy estimate closely related to, but not identical with, the conceptual ITFR (note that exact WVD “optimality” holds for linear FM components as discussed in Section II-B). Regardless, these residual WVD “artificial” auto terms are substantially mitigated in the RIFT through the employment of the entropic weighting scheme, CFWT wavelet basis, and the positivity-constrained LR–TV process, which suppress sign-changing (positive and negative) oscillatory artefacts (as observed in the synthetic signal results).
In terms of computational complexity, let $V$ be the number of time samples, $U$ the number of frequency bins in the T–F grid, and $N_{k}=\sum_{n=1}^{N}M(\sigma_{n})$, the total number of CFWT kernels. Given also that the LR deconvolution process runs for $K_{\mathrm{LR}}$ iterations, the RIFT complexity can be derived to be $\approx\mathcal{O}\!\left((N_{k}+K_{\mathrm{LR}})\,U\,V\log(UV)\right)$. Given that the WVD costs $\mathcal{O}\!\big(V^{2}+V\,U\log U\big)$, with Cohen’s class methods requiring an additional $\mathcal{O}(U\,V\log(UV))$, for $U\!\approx\!V$, the RIFT, WVD, and Cohen’s class methods are all $\mathcal{O}(UV\log U)$ up to constants. The RIFT trades factors $N_{k}$ and $K_{\mathrm{LR}}$ against WVD’s $V^{2}$ and Cohen’s extra 2-D smoothing. For long signals ($V\!\gg\!U$), the $V^{2}$ term dominates WVD-based methods, so for small-$N_{k}$, the RIFT can be cheaper; for frequency-heavy grids ($U\!\gg\!V$), costs are comparable up to logarithmic factors. For further details, see Section J in the SM.
@@ IV Results
Let $R=|\text{T--F Representation}|$, and $C=\text{ITFR}_{ref}$. The reference ITFR, $\text{ITFR}_{ref}$, is rendered as per the ITFR model in Section II-A, rasterised with Xiaolin Wu’s antialiased line algorithm [29] and lightly convolved with an isotropic Gaussian ($\sigma_{I}=1.5$px) to form a tolerance tube to ensure metric stability against discretisation and sub-pixel jitter; note that ablating $\sigma_{I}$ confirms metric ranking insensitivity to this parameter (see Table II in the SM). Let $P=R/\sum_{t,f}R$ and $Q=C/\sum_{t,f}C$. The following metrics are reported: the Bhattacharyya overlap quantifies global probabilistic overlap between $P$ and $Q$: $\mathrm{BC}=\sum_{t,f}\sqrt{P\,Q}$ [56] (higher is better). The Jensen–Shannon divergence measures the symmetric, finite KL discrepancy between $P$ and $Q$: $\mathrm{JS}=\tfrac{1}{2}\mathrm{KL}(P\|M)+\tfrac{1}{2}\mathrm{KL}(Q\|M)$, $M=\tfrac{1}{2}(P+Q)$ [57] (lower is better). The Ridge Energy Ratio captures on-ridge concentration as the fraction of raw energy inside the tube weights $W=C/\max C$: $\mathrm{RER}=\sum_{t,f}P\,W$  [58, 59] (higher is better). The metrics are reported for varying additive white Gaussian noise (AWGN) signal-to-noise ratios (SNRs) to test noise tolerance: 10,5,0,5-10,-5,0,5, and $\infty\,\text{dB}$ (noise-free signal). Results for SNRs 5-5 and $\infty\,\text{dB}$ are shown for all evaluated methods in Fig. 4, and metrics for all SNRs are reported in table I; per-example metrics and remaining SNR figures can be found in Tables III-IV and Figs. 10-11 respectively in the SM. To ensure readability at publication scale, we apply a small isotropic Gaussian anti-aliasing prefilter ($\sigma=1$px) for visualisation only, uniformly across methods and SNRs; all quantitative metrics are computed from the unfiltered TFRs.
@@ V Conclusion
TABLE I: T–F Evaluation metrics across SNRs: -10, -5, 0, +5, and $\infty$ (dB). Bold indicates the best value per row; arrows denote the preferred direction (↑ = higher, ↓ = lower). Cells report the mean across the two examples (equal weight); per-example results are provided in Fig. III and IV in the SM.
@@
B Demonstration of CFWT T–F trajectory alignment
Consider the following paired chirp signal:
$z_{3}(t)$ $=e^{j\phi_{1}(t)}+e^{j\phi_{2}(t)}$ (77)
where:
$\phi_{1}(t)=2\pi\left(50t+\frac{50}{3}t^{2}\right),\qquad\phi_{2}(t)=2\pi\left(60t+\frac{50}{3}t^{2}\right).$ (78)
The WVD of $z_{3}(t)$ can be derived as:
$WVD_{z_{3}}\left(f,t\right)=$ $\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\delta\left(f-\left[50+\dfrac{100}{3}\,t\right]\right)+\delta\left(f-\left[60+\dfrac{100}{3}\,t\right]\right)}}_{\text{``Auto-terms'' (desirable terms)}}$
$+\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2\cos(20\pi t)\,\delta\left(f-\left[55+\dfrac{100}{3}\,t\right]\right)}}_{\text{Cross-term}}.$ (79)
Figure 5 presents the derived WVD, the standard CWT applied for small and large standard deviations, and the generalised CFWT presented in Eq. (15) with $\theta$ appropriately set. As observed in the plot, choosing a suitable $\theta$ value aligns the resulting T–F kernel with the signal trajectories and consequently the cross-term. Thus, convolution with this kernel (providing the standard deviation is sufficiently high) substantially reduces the cross-term and isolates the auto-terms (desirable terms) with high resolution.
Figure 5: Comparison between the (a) WVD (red +, blue -), the standard CWT applied for (b) small and (c) large standard deviations, and the (d) generalised CFWT presented in Eq. (15) aligned with the signal trajectory, for signal $z_{3}(t)$.
C Derivation of the Cohen’s class time–frequency convolution kernel
@@
$=\frac{1}{2\sqrt{\pi^{3}\sigma^{2}}}e^{-\left[\frac{t^{2}}{\sigma^{2}}+\sigma^{2}\left(\omega-\tan{\left(\theta\right)}\,t\right)^{2}\right]}2\sqrt{\pi\sigma^{2}}.$
$=\frac{1}{\pi}e^{-\left[\frac{t^{2}}{\sigma^{2}}+\sigma^{2}\left(\omega-\tan{\left(\theta\right)}\,t\right)^{2}\right]}.$
where:
D Derivation of the Cohen’s class kernel principal axis angular and standard deviation offset
@@
with $a,b,c$ taken from Eq. (81). Thus, the variance (eigenvalue) along the principal axis, $\lambda_{1}$, is
@@
E Visualisation of the effect of the Wavelet function on the ITFR
The general T–F characteristics of the presented wavelet function can be visualised by considering signal events in the extremes of resolution: a delta function and an infinitely long sinusoid. Consider the Fourier Transform of the wavelet function for $\theta=0$:
$\bar{W}_{\omega}\left(\phi\right)=\mathcal{F}\left\{W_{\omega}\left(t\right)\right\}$
$=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}\frac{1}{\sqrt[4]{\pi\sigma^{2}}}e^{-\frac{1}{2\sigma^{2}}\left(t^{2}-2i\omega\sigma^{2}t\right)}e^{-i\phi t}\,dt$
$=\int_{\mathbb{R}}\frac{1}{\sqrt[4]{4\pi^{3}\sigma^{2}}}e^{-\frac{1}{2\sigma^{2}}\left(t+i\left(\phi-\omega\right)\sigma^{2}\right)^{2}-\frac{\left(\phi-\omega\right)^{2}\sigma^{2}}{2}}\,dt$
$=\sqrt[4]{\frac{\sigma^{2}}{\pi}}e^{-\frac{\left(\phi-\omega\right)^{2}\sigma^{2}}{2}}\int_{\mathbb{R}}I_{G}\left(t\right)\,dt,$ (97)
where:
$I_{G}\left(t\right)$ $=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{1}{2\sigma^{2}}\left(t+i\left(\phi-\omega\right)\sigma^{2}\right)^{2}}$
$=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{1}{2\sigma^{2}}\left(t+i\mu\right)^{2}}$ (98)
To evaluate the integral, a suitable complex contour, $\Gamma$ must be constructed and Cauchy’s theorem employed:
$\oint_{\Gamma}I_{z}\left(z\right)=0,$ (99)
where:
$I_{z}\left(z\right)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{1}{2\sigma^{2}}z^{2}}.$ (100)
Thus:
$\int_{-a}^{a}I_{z}\left(x\right)\,dx+\int_{0}^{\mu}I_{z}\left(a+i\mu\right)\,dy$
$\hskip 14.22636pt+\int_{a}^{-a}I_{z}\left(x+i\mu\right)\,dx+\int_{\mu}^{0}I_{z}\left(-a+i\mu\right)\,dy=0.$ (101)
Therefore:
$-\int_{a}^{-a}I_{z}\left(x+i\mu\right)\,dx$ $=\int_{-a}^{a}I_{G}\left(x\right)\,dx$
$=\int_{-a}^{a}I_{z}\left(x\right)\,dx+\int_{0}^{\mu}I_{z}\left(a+i\mu\right)\,dy$
$\hskip 14.22636pt+\int_{\mu}^{0}I_{z}\left(-a+i\mu\right)\,dy.$ (102)
Taking limits:
$\lim_{a\to\infty}\left\{-\int_{a}^{-a}I_{z}\left(x+i\mu\right)\,dx\right\}$ $=\int_{\mathbb{R}}I_{z}\left(x\right)\,dx=1.$ (103)
Therefore, $\bar{W}_{\omega}\left(\phi\right)$ can be expressed as:
$\bar{W}_{\omega}\left(\phi\right)=\sqrt[4]{\frac{\sigma^{2}}{\pi}}e^{-\frac{\left(\phi-\omega\right)^{2}\sigma^{2}}{2}}.$ (104)
Note that, as expected from the energy principle in (32):
$E_{W}$ $=\int_{\mathbb{R}}\left|\bar{W}_{\omega}\left(\phi\right)\right|^{2}\,d\omega$
$=\int_{\mathbb{R}}\sqrt{\frac{\sigma^{2}}{\pi}}e^{-\left(\phi-\omega\right)^{2}\sigma^{2}}\,d\omega$
$=\sqrt{\frac{\sigma^{2}}{\pi}}\int_{\mathbb{R}}e^{-u^{2}\sigma^{2}}\,du$
$=1.$ (105)
Let $y\left(t\right)=x\left(t\right)*W_{\omega}\left(t\right)$. Thus, for an input frequency, $x(t)=e^{j\omega_{0}t}$ ($\omega_{0}>0$), $y\left(t\right)$ can be evaluated in the frequency domain:
$\bar{Y}\left(\phi\right)$ $=\mathcal{F}\left\{y(t)\right\}$
$=\mathcal{F}\left\{x\left(t\right)*W_{\omega}\left(t\right)\right\}$
$=\sqrt{2\pi}\mathcal{F}\left\{e^{j\omega_{0}t}\right\}\cdot\mathcal{F}\left\{W_{\omega}\left(t\right)\right\}$
$=2\pi\delta\left(\phi-\omega_{0}\right)\sqrt[4]{\frac{\sigma^{2}}{\pi}}e^{-\frac{\left(\phi-\omega\right)^{2}\sigma^{2}}{2}}$
Thus, taking the magnitude squared of the inverse Fourier Transform:
$\left|y\left(\omega,\,t\right)\right|^{2}$ $=\left|\mathcal{F}^{-1}\left\{\bar{Y}\left(\omega_{0}\right)\right\}\right|^{2}$
$=\left|\frac{1}{2\pi}\int_{\mathbb{R}}2\pi\delta\left(\phi-\omega_{0}\right)\sqrt[4]{\frac{\sigma^{2}}{\pi}}e^{-\frac{\left(\phi-\omega\right)^{2}\sigma^{2}}{2}}e^{i\phi t}\,d\phi\right|^{2}$
$=\sqrt{\frac{\sigma^{2}}{\pi}}e^{-\left(\omega_{0}-\omega\right)^{2}\sigma^{2}}.$
Thus, the process equivalently smooths the frequency precision of the input signal (of frequency $\omega_{0}$) by a Gaussian with a standard deviation equal to the inverse of the wavelet parameter $\sigma$ (divided by $\sqrt{2}$). Likewise, taking an input signal $x(t)=\delta(t)$ reveals the temporal behaviour of $f\left(\omega,\,t\right)$:
$\left|y\left(\omega,\,t\right)\right|^{2}$ $=\left|\left[\delta*W_{\omega}\right]\left(t\right)\right|^{2}$
$=\left|\int_{\mathbb{R}}\delta(\tau)W_{\omega}\left(t-\tau\right)\,d\tau\right|^{2}$
$=\left|W_{\omega}\left(t\right)\right|^{2}$
$=\frac{1}{\sqrt{\pi\sigma^{2}}}e^{-\frac{1}{\sigma^{2}}t^{2}}\left|e^{i\omega t}\right|$
$=\frac{1}{\sqrt{\pi\sigma^{2}}}e^{-\frac{1}{\sigma^{2}}t^{2}}.$ (106)
Thus, the process equivalently smooths the temporal precision of the input signal by a Gaussian with standard deviation equal to that of the wavelet window function ($\sigma$) (multiplied by $\sqrt{2}$). Note that evaluating the WVD for the oscillatory signal yields:
$WVD\left\{Ae^{j\omega_{0}t}\right\}\left(\omega,t\right)$ $=\frac{1}{2\pi}\int_{\mathbb{R}}e^{j\omega_{0}\left(t+\frac{\tau}{2}\right)}e^{-j\omega_{0}\left(t-\frac{\tau}{2}\right)}e^{-j\omega\tau}\,d\tau$
$=\frac{1}{2\pi}\int_{\mathbb{R}}e^{j2\pi\left(f_{0}-\frac{\omega}{2\pi}\right)\tau}\,d\tau$
$=\frac{1}{2\pi}\delta\left(f_{0}-\frac{\omega}{2\pi}\right),$ (107)
which corresponds to perfect frequency resolution. Likewise for the impulse:
$WVD\left\{\delta(t)\right\}\left(\omega,t\right)$ $=\frac{1}{2\pi}\int_{\mathbb{R}}\delta\left(t+\frac{\tau}{2}\right)\delta\left(t-\frac{\tau}{2}\right)e^{-j\omega\tau}\,d\tau$
$=\frac{1}{2\pi}\delta(t),$ (108)
which corresponds to perfect time resolution.
F Derivation of the Reconstructive Ideal Fractional Transform

### v3

@@ III-A Proposed Wavelet Transform
$\theta$ is the intended angle of the instantaneous angular frequency relative to the horizontal axis (time axis) in the T–F plane, and $\sigma$ is the intended standard deviation of the major axis of the corresponding Cohen’s class kernel. As explored previously in the overview (Section I-B), the employment of Short-time Fractional Fourier (or wavelet) Transforms is especially desirable in the context of time-varying signal curvature compared to their standard counterparts (for instance Short-Time Fourier Transforms or Wavelet Transforms), given their ability to align with signal trajectories and thus minimise cross term interference appropriately [11]. Employing the result of Eq. (14), the exact Cohen’s class smoothing kernel, $\bar{\Pi}_{\sigma,\theta}\left(\omega,\,t\right)$, resulting from the wavelet convolutional process can be obtained as:
@@
For the derivation, see the SM, Section B. Note that due to the skewed form of the kernel in Eq. (19), the intended angle and standard deviation of the principal axis of the kernels may not necessarily coincide with the actual values. The required mapping between the intended pair $(\sigma,\theta)$ (principal-axis angle and standard deviation) to the implemented parameters $(\sigma_{0},\kappa)$ is given by:
@@
where $K_{\sigma}\!\left(\kappa\right)$ is the axis switching function provided in Eq. (90) in the SM. Thus, $\kappa(\sigma,\theta)=\varphi^{-1}(\sigma,\theta)$, the inverse function with respect to $\theta$. Note that in the limiting case $\sigma\to\infty$, we obtain $\varphi(\sigma,\theta)\to\theta$ as expected. Equivalently, the mapping $(\sigma,\theta)\mapsto(\sigma_{0},\kappa)$ can be obtained by performing covariance matching with the intended Cohen’s class kernel (introduced in Eq. (25)); further details and derivations are given in the SM, Sections C1 and C2.
@@ III-B Reconstructing the ITFR
See SM, Section D, for the full derivation. However, it is known that the solution for $\hat{\text{ITFR}}$ needs to be strictly non-negative from Eq. (33). Thus, we propose a nonnegativity-constrained solution to these normal equations (Eq. (98)) via Lucy–Richardson Deconvolution with total variation regularisation (LR-TV) [48]. TV regularisation, implemented via the multiplicative divergence term in Eq. (40) below, penalises the sum over T–F pixels of the Euclidean ($\ell_{2}$) norm of the local T–F gradient (isotropic TV). TV suppresses isolated pixel-scale speckle and deconvolution ringing while preserving edge-like ridges [48]. This standard LR-TV form [48] yields the following iterative update for the RIFT estimate of the ITFR, with respect to a spatially varying PSF, $\mathbf{\Psi}$:
@@
To efficiently compute the spatially varying Lucy–Richardson (LR) process, the overlap-save method can be employed [52, 53]. This approach involves performing parallel LR deconvolutions on $L\times H$ partitioned blocks of the 2D array. The results are then “stitched” together (with appropriate padding) to approximate the spatially varying process while maintaining computational efficiency. Full implementation details are provided in Section E in the SM. Likewise, a breakdown of the global RIFT process is provided in Algorithm 1.
@@ III-D Entropic Filtering for Cross-term Suppression
It is recommended that $w_{u,v}$ is a bivariate Gaussian, with kernel variances selected to optimise performance for the desired application. For instance, for an instrumental or speech extract with known linearly spaced harmonics, the $\omega$ axis variance can be set high to enforce harmonic trajectory alignment in the presence of noise, otherwise, the kernel can be set to an isotropic Gaussian of variance greater than that of the derived variance of the isotropic Gaussian kernel in Eq. (45) to ensure valid entropic weightings.
@@ III-E Fundamental Properties of the RIFT
Note that, by construction, the RIFT output, $\widehat{I}(\omega,t)$, is positive ($\widehat{I}(\omega,t)\!\geq\!0$), time–shift and frequency–shift equivariant up to discretisation ($\widehat{I}_{x_{\tau}}(\omega,t)=\widehat{I}_{x}(\omega,t-\tau)$ for $x_{\tau}(t)=x(t-\tau)$ and $\widehat{I}_{x_{\nu}}(\omega,t)=\widehat{I}_{x}(\omega-\nu,t)$ for $x_{\nu}(t)=x(t)\,e^{j2\pi\nu t}$). See Section F in the SM for more details. It is also non-linear, specifically quadratic, due to the magnitude–squared of the CFWT coefficients (corresponding to energy). Likewise, by construction, the RIFT estimates the squared modulus of the ideal time–frequency representation (no phase information) and is thus non-invertible. Also, note that whilst the RIFT substantially suppresses cross-terms, it does not guarantee their elimination in all cases. Likewise, the target is a cross-term-free, WVD-resolution energy estimate closely related to, but not identical with, the conceptual ITFR (note that exact WVD “optimality” holds for linear FM components as discussed in Section II-B). Regardless, these residual WVD “artificial” auto terms are substantially mitigated in the RIFT through the employment of the entropic weighting scheme, CFWT wavelet basis, and the positivity-constrained LR–TV process, which suppress sign-changing (positive and negative) oscillatory artefacts (as observed in the synthetic signal results).
In terms of computational complexity, let $V$ be the number of time samples, $U$ the number of frequency bins in the T–F grid, and $N_{k}=\sum_{n=1}^{N}M(\sigma_{n})$, the total number of CFWT kernels. Given also that the LR deconvolution process runs for $K_{\mathrm{LR}}$ iterations, the RIFT complexity can be derived to be $\approx\mathcal{O}\!\left((N_{k}+K_{\mathrm{LR}})\,U\,V\log(UV)\right)$. Given that the WVD costs $\mathcal{O}\!\big(V^{2}+V\,U\log U\big)$, with Cohen’s class methods requiring an additional $\mathcal{O}(U\,V\log(UV))$, for $U\!\approx\!V$, the RIFT, WVD, and Cohen’s class methods are all $\mathcal{O}(UV\log U)$ up to constants. The RIFT trades factors $N_{k}$ and $K_{\mathrm{LR}}$ against WVD’s $V^{2}$ and Cohen’s extra 2-D smoothing. For long signals ($V\!\gg\!U$), the $V^{2}$ term dominates WVD-based methods, so for small-$N_{k}$, the RIFT can be cheaper; for frequency-heavy grids ($U\!\gg\!V$), costs are comparable up to logarithmic factors.
@@ IV Results
Let $R=|\text{T--F Representation}|$, and $C=\text{ITFR}_{ref}$. The reference ITFR, $\text{ITFR}_{ref}$, is rendered as per the ITFR model in Section II-A, rasterised with Xiaolin Wu’s antialiased line algorithm [29] and lightly convolved with an isotropic Gaussian ($\sigma_{I}=1.5$px) to form a tolerance tube to ensure metric stability against discretisation and sub-pixel jitter; note that ablating $\sigma_{I}$ confirms metric ranking insensitivity to this parameter (see Table II in the SM). Let $P=R/\sum_{t,f}R$ and $Q=C/\sum_{t,f}C$. The following metrics are reported: the Bhattacharyya overlap quantifies global probabilistic overlap between $P$ and $Q$: $\mathrm{BC}=\sum_{t,f}\sqrt{P\,Q}$ [56] (higher is better). The Jensen–Shannon divergence measures the symmetric, finite KL discrepancy between $P$ and $Q$: $\mathrm{JS}=\tfrac{1}{2}\mathrm{KL}(P\|M)+\tfrac{1}{2}\mathrm{KL}(Q\|M)$, $M=\tfrac{1}{2}(P+Q)$ [57] (lower is better). The Ridge Energy Ratio captures on-ridge concentration as the fraction of raw energy inside the tube weights $W=C/\max C$: $\mathrm{RER}=\sum_{t,f}P\,W$  [58, 59] (higher is better). The metrics are reported for varying additive white Gaussian noise (AWGN) signal-to-noise ratios (SNRs) to test noise tolerance: 10,5,0,5-10,-5,0,5, and $\infty\,\text{dB}$ (noise-free signal). Results for SNRs 5-5 and $\infty\,\text{dB}$ are shown for all evaluated methods in Fig. 4, and metrics for all SNRs are reported in Table I. To ensure readability at publication scale, we apply a small isotropic Gaussian anti-aliasing prefilter ($\sigma=1$px) for visualisation only, uniformly across methods and SNRs; all quantitative metrics are computed from the unfiltered TFRs.
@@ V Conclusion
TABLE I: T–F Evaluation metrics across SNRs: -10, -5, 0, +5, and $\infty$ (dB). Bold indicates the best value per row; arrows denote the preferred direction (↑ = higher, ↓ = lower). Cells report the mean across the two examples (equal weight).
@@
B Derivation of the Cohen’s class time–frequency convolution kernel
@@
$=\frac{1}{2\sqrt{\pi^{3}\sigma^{2}}}e^{-\left[\frac{t^{2}}{\sigma^{2}}+\sigma^{2}\left(\omega-\tan{\left(\theta\right)}\,t\right)^{2}\right]}2\sqrt{\pi\sigma^{2}}$
$=\frac{1}{\pi}e^{-\left[\frac{t^{2}}{\sigma^{2}}+\sigma^{2}\left(\omega-\tan{\left(\theta\right)}\,t\right)^{2}\right]}.$ (77)
C Derivation of the Cohen’s class kernel principal axis angular and standard deviation offset
@@
with $a,b,c$ taken from Eq. (20) from the main paper. Thus, the variance (eigenvalue) along the principal axis, $\lambda_{1}$, is
@@
D Derivation of the Reconstructive Ideal Fractional Transform